# Differentiable Fuzzy Cosmic-Web for Field Level Inference

### v1

@@ 2.8 Differentiable Fuzzy Cosmic-Web Classification
In the top-down structure formation scenario proposed by Zel’dovich (1970), cosmic structures form via anisotropic collapse governed by the tidal field tensor. The process unfolds along the tensor’s eigenvectors, with collapse first occurring along the direction associated with the largest eigenvalue–a phenomenon known as pancake formation. The term cosmic-web was coined by Bond et al. (1996) to describe the large-scale structure of the Universe, characterized by the emergence of distinct morphological components–such as filaments, walls, and nodes–arising from anisotropic gravitational collapse. This interpretation was supported by shape parameters like ellipticity and prolateness, derived from the eigenvalues of the deformation tensor, and linked to the Zel’dovich approximation. A more systematic and operational classification of the cosmic-web was later proposed by Hahn et al. (2007), who used the eigenvalues of the tidal (equal to the velocity shear within the Zel’dovich approximation) tensor to categorize regions of space into knots, filaments, sheets, and voids, depending on the number of collapsing directions. This method provided a quantitative tool to identify web environments in cosmological simulations.
While widely used, cosmic-web classification schemes remained largely phenomenological, lacking a firm quantitative foundation. A more rigorous interpretation began to emerge with the connection to the framework of nonlocal bias, which provided a physically motivated context for understanding how large-scale structure influences halo and galaxy formation–an approach we adopt in this work (see Appendix B).
Let us consider a cosmic-web classification scheme based on the tidal field tensor, $\partial_{j}\partial_{k}\Phi$ , where $\Phi=\nabla^{-2}\delta$ is the gravitational potential. Let $\lambda_{i}^{(1)}<\lambda_{i}^{(2)}<\lambda_{i}^{(3)}$ be the eigenvalues of the tidal field tensor at the $i$th cell. The cosmic-web classification at the $i$th cell can be compactly expressed as
@@
where $h$ is the Heaviside step function and $\lambda_{\text{th}}$ is a free parameter. In this work we set $\lambda_{\text{th}}=0.05$ (see Forero–Romero et al., 2009, for the introduction of a threshold in the cosmic-web classification). For recent studies regarding the choice of $\lambda_{\text{th}}$, see Coloma-Nadal et al. (2024); Olex et al. (2025). $W_{i}$ counts the number of eigenvalues above the threshold; the four resulting values 0, 1, 2, 3 correspond to voids, sheets, filaments, and knots, respectively. This cosmic-web definition is problematic if introduced in a forward model that aims to be differentiable with respect to $\boldsymbol{\nu}$. This is because the cosmic-web classification mesh $\boldsymbol{W}$ depends on $\boldsymbol{\nu}$ through the potential of $\boldsymbol{\delta}=\mathcal{M}_{\Psi}\circ\mathcal{M}_{\delta}(\boldsymbol{\nu})$, but it does so in a manifestly non-differentiable way due to the presence of the Heaviside step function. As a consequence, the gradients $\partial W_{i}/\partial q_{j}$ are ill-defined, which precludes the use of gradient-based sampling methods. To address this we rely on a fuzzy cosmic-web classification defined through the sigmoid weights, with
$w^{(n)}_{i}=\sigma[s(\lambda^{(n)}_{i}-\lambda_{\text{th}})]$ (12)
@@
where $\sigma(x)=(1+\mathrm{e}^{-x})^{-1}$, and $s$ is the steepness parameter controlling how sharply $w^{(n)}_{i}$ transitions around $\lambda_{\text{th}}$. Because $\sigma$ is strictly increasing and the eigenvalues are sorted, the weights inherit the order $w^{(1)}>w^{(2)}>w^{(3)}$. We convert these cumulative exceedance probabilities into mutually exclusive “fuzzy” memberships for knots (K), filaments (F), sheets (S) and voids (V):
$p^{(\text{K})}_{i}$ $=w^{(3)}_{i}\ ,$ (13)
$p^{(\text{F})}_{i}$ $=w^{(2)}_{i}-w^{(3)}_{i}\ ,$ (14)
$p^{(\text{S})}_{i}$ $=w^{(1)}_{i}-w^{(2)}_{i}\ ,$ (15)
$p^{(\text{V})}_{i}$ $=1-w^{(1)}_{i}\ ,$ (16)
@@
so that they represent the probabilities of having exactly three, two, one, or zero eigenvalues above the threshold and, by construction, $p^{(\mathrm{V})}_{i}+p^{(\mathrm{S})}_{i}+p^{(\mathrm{F})}_{i}+p^{(\mathrm{K})}_{i}=1$. By introducing this smooth classification scheme the forward model remains differentiable. Regarding the steepness parameter $s$, a larger value makes the fuzzy classification closer to the original classification, but it also risks causing numerical instabilities in the gradients. We therefore choose the largest $s$ that preserves stable gradient computations. In Figure 2 we show a comparison of the “sharp” and smooth cosmic-web classification.
@@
We now consider a cosmic-web dependent bias model where each of the bias parameters from the local bias model depend on the cosmic-web classification of the voxel. That is, given a bias parameter $b$, it now takes different values in each of the cosmic-web regions: $b^{(\text{K})},b^{(\text{F})},b^{(\text{S})},b^{(\text{V})}$. To ensure differentiability, we use the fuzzy cosmic-web classification, and the bias parameters at voxel $i$ takes the value
$b_{i}=b^{(\text{K})}p^{(\text{K})}_{i}+b^{(\text{F})}p^{(\text{F})}_{i}+b^{(\text{S})}p^{(\text{S})}_{i}+b^{(\text{V})}p^{(\text{V})}_{i}\ .$ (17)
2.9 Redshift-Space Distortions Modelling
@@
Redshift-space distortions (RSDs) arise when the observed position of a tracer deviates from its Hubble-flow distance due to peculiar velocities, making them sensitive probes of the growth rate of structure formation (see Kaiser, 1987; Hamilton, 1998). Redshift-space distortions are governed by two distinct types of peculiar motions: coherent large-scale flows, which arise from gravitational infall toward overdense regions, and small-scale random motions within quasi-virialized structures. For a variety of ways of modelling coherent flows we refer to Kitaura et al. (2012a) and references therein (see also Wang et al., 2012). There are several strategies to incorporate RSDs within a Bayesian framework (we choose the third):
@@
    Iterative sampling of the redshift-to-real-space mapping:
    This approach iteratively reconstructs the real-space distribution from redshift-space observations using Gibbs sampling (Kitaura & Enßlin, 2008; Kitaura et al., 2012c, 2016a). It is inspired by earlier iterative RSD correction methods (see Yahil et al., 1991; Monaco & Efstathiou, 1999).
@@
    Forward modelling of RSDs at the tracer level:
    Originally introduced by Kitaura et al. (2012b), this method evolves tracers from sampled initial conditions to redshift space, where their final positions are compared with observed galaxy distributions to constrain the primordial density field. Small-scale virial motions can be addressed in two ways: either by collapsing elliptical groups of galaxies–as proposed by Tegmark et al. (2004)–and modelling only coherent flows, as in the original study; or by retaining Fingers-of-God features and modelling them with a stochastic velocity dispersion component, as done by Heß et al. (2013).
@@
    Forward modelling of RSDs at the dark matter field level: In this work, we follow the approach of mapping the matter field directly to redshift-space in a fully differentiable Bayesian inference framework. We set the observer at the center of the cubical volume and compute the RSD effects accordingly along line-of-sights. See Bos et al. (2019) for a first implementation of such a method and calculation details.
We note that the modelling of RSDs for biased tracers involving nonlinear transformations–such as cosmic voids and Lyman-$\alpha$ forests–requires the inclusion of velocity bias. For the Lyman-$\alpha$ forest, RSD affects the optical depth $\tau$, but the observable is the transmitted flux $F=\exp(-\tau)$, introducing a nontrivial mapping. Accurate modelling in these cases must account for velocity bias effects McDonald et al. (2000); Seljak (2012); Sinigaglia et al. (2022, 2024b). Similarly, cosmic voids are identified from the galaxy distribution in redshift space, which effectively applies a nonlinear and nonlocal transformation Chuang et al. (2017).
@@ 3 Validation of Field-Level Inference with Fuzzy Cosmic-Web Bias
To validate the BRIDGE code, we consider three numerical tests with different resolutions and bias models: TEST1, TEST2, TEST3 (see Tab. 1). Results of TEST2 and TEST3 are presented in Appendix C. The numerical tests consist of the following steps:
@@
    Generate a ground-truth number counts catalog of biased tracers from an initial Gaussian density field with the same forward model employed during inference. All runs are performed to redshift $z=0$ with a particular random seed and specific bias parameters as listed in Tables 2 and 3.
@@
    Field-level inference of the posterior describing the white-noise–and bias parameters for cases TEST1 and TEST2–applying the BRIDGE code.
@@
    Quality assessment of the reconstructions is performed by evaluating the recovery of both the initial and final conditions using two-point and three-point statistical analyses.
Table 1: Setting of the different numerical cases. In all cases we use a mesh of $128^{3}$ voxels with cubical volumes of distinct side lengths $L$ and resolutions ${\Delta}L=L/N$. We consider the Nyquist frequency $k_{\rm nyq}=\pi/\Delta L$ and the isotropic Nyquist frequency $k^{\rm iso}_{\rm nyq}=k_{\rm nyq}/\sqrt{3}$ of the sphere contained in the cubical Fourier-space mesh. The output redshift is $z=0$. There are 16 combinations of $\Phi$- and $\delta$-web regions: $j,k\in[1,2,3,4]$.
@@
$\Delta L$ [$h^{-1}\,\mathrm{Mpc}$] 10 5 8
$L$ [$h^{-1}\,\mathrm{Mpc}$] 1280 640 1024
$k^{\rm iso}_{\rm nyq}$ [$h\,\mathrm{Mpc}^{-1}$] 0.31 0.63 0.39
$k_{\rm nyq}$ [$h\,\mathrm{Mpc}^{-1}$] 0.54 1.10 0.68
nonlocal bias $\Phi$-webj $\Phi$-webj $\Phi\delta$-webjk
@@
Table 2: Values of bias parameters by cosmic-web region of the reference for the TEST1 & TEST2 cases (8 bias parameters)
@@
Table 3: Values of bias parameters by combined cosmic-web regions for the TEST3 case (64 bias parameters), showing subdivisions of $\delta$-web types within each $\Phi$-web classification (see Appendix B and Coloma-Nadal et al. (2024)).
@@ 3.1 Motivation of the Numerical Tests
The numerical tests were chosen to demonstrate a number of scientific relevant cases:
@@
  • TEST1 & TEST2: The nonlocal bias description used here is based on the $\Phi$-web classification. Balaguera-Antolínez et al. (2018) demonstrated that this model can accurately reproduce key statistical properties of the halo distribution, including one-point PDFs, two-point power spectra at percent-level accuracy, and bispectra with reasonable agreement–in general within 15% (see also Balaguera-Antolínez et al., 2020; Pellejero-Ibañez et al., 2020; Balaguera-Antolínez et al., 2023, confirming these results for different halo resolutions). These studies confirm the $\Phi$-web as a physically meaningful and effective framework for modeling nonlocal bias within a reasonable degree of accuracy through non-parametric approaches. However, for field-level inference applications–such as the one developed here a parametric bias model is required to enable efficient sampling and marginalization within a Bayesian framework. The same nonlocal bias treatment incorporating an equivalent nonlinear bias prescription to the one considered here (see Eq. 7)–was employed by Sinigaglia et al. (2024b, a) to model the Lyman-$\alpha$ forest. Their work demonstrated excellent agreement with high-resolution hydrodynamical simulations, achieving power spectrum accuracy within 5% up to $k\sim 1$ $h\,\mathrm{Mpc}^{-1}$, along with consistent bispectrum predictions. In both cases–haloes and Lyman-$\alpha$ forests–the underlying structure formation model was ALPT. For both TEST1 & TEST2 cases we jointly sample the bias parameters starting with an initial guess randomly sampled from the prior distribution (see Eq. 4).
@@
  • TEST1 focuses on a resolution of 10 $h^{-1}\,\mathrm{Mpc}$, corresponding to an isotropic Nyquist frequency of $k\simeq 0.3$ $h\,\mathrm{Mpc}^{-1}$. This resolution surpasses the typical scale used in most cosmological analyses, which are commonly limited to $k<0.2$ $h\,\mathrm{Mpc}^{-1}$(see, e.g., Ivanov et al., 2025). Also the ideal resolution at which BAO reconstruction techniques are applied is slightly lower–15 $h^{-1}\,\mathrm{Mpc}$(see, e.g., Paillas et al., 2025).
@@
  • TEST2 focuses on a resolution of 5 $h^{-1}\,\mathrm{Mpc}$ with a Nyquist frequency of $k\simeq 1$ $h\,\mathrm{Mpc}^{-1}$. This resolution has been reported to be high enough to produce accurate galaxy catalogs combined with a subgrid model based on ALPT (see Forero Sánchez et al., 2024).
@@
  • TEST3: This case study incorporates the full HICOBIAN nonlocal bias model, which has been shown to significantly enhance accuracy in both power spectra and bispectra, achieving full compatibility with $N$-body-based catalogs across a wide range of scales. While Coloma-Nadal et al. (2024) demonstrated that this model remains valid at mesh resolutions below 4 $h^{-1}\,\mathrm{Mpc}$, in this study we restrict our analysis to a coarser resolution of 8 $h^{-1}\,\mathrm{Mpc}$. This choice is motivated by the fact that the corresponding isotropic Nyquist frequency, $k\simeq 0.4$ $h\,\mathrm{Mpc}^{-1}$, marks the scale beyond which cosmological information is effectively saturated (Hahn & Villaescusa-Navarro, 2021). For computational reasons, we keep the bias parameters fixed in this case, assuming they are known. This assumption is reasonable if the parameters can be extracted from high-fidelity reference catalogs, as was done for the PATCHY BOSS mocks, where the MULTIDARK reference catalog was designed to reproduce the observed galaxy distribution with Halo Abundance Matching (Rodríguez-Torres et al., 2016). Even more sophisticated reference catalogs are now being developed to reproduce the clustering statistics of the observed DESI catalogs (DESI Collaboration et al., 2025), from which the parameters of the HICOBIAN model can be accurately extracted (Favole, G. et al., 2025).

### v3

@@ 2.8 Differentiable fuzzy cosmic-web classification
In the top-down structure-formation scenario proposed by Zel’dovich (1970), cosmic structures form via anisotropic collapse governed by the tidal field tensor. The process unfolds along the tensor’s eigenvectors, with collapse first occurring along the direction associated with the largest eigenvalue – a phenomenon known as pancake formation. The term cosmic web was coined by Bond et al. (1996) to describe the large-scale structure of the Universe, characterised by the emergence of distinct morphological components such as filaments, walls, and nodes arising from anisotropic gravitational collapse. This interpretation was supported by shape parameters such as ellipticity and prolateness, derived from the eigenvalues of the deformation tensor and linked to the Zel’dovich approximation. A more systematic and operational classification of the cosmic web was later proposed by Hahn et al. (2007), the authors of which used the eigenvalues of the tidal (equal to the velocity shear within the Zel’dovich approximation) tensor to categorise regions of space into knots, filaments, sheets, and voids, depending on the number of collapsing directions. This method provided a quantitative tool to identify web environments in cosmological simulations.
While widely used, cosmic-web classification schemes remained largely phenomenological, lacking a firm quantitative foundation. A more rigorous interpretation began to emerge with the connection to the framework of non-local bias, which provided a physically motivated context for understanding how large-scale structure influences halo and galaxy formation – an approach we adopted in this work (see Appendix B).
We now consider a cosmic-web classification scheme based on the tidal field tensor, $\partial_{j}\partial_{k}\Phi$ , where $\Phi=\nabla^{-2}\delta$ is the gravitational potential. Let $\lambda_{i}^{(1)}<\lambda_{i}^{(2)}<\lambda_{i}^{(3)}$ be the eigenvalues of the tidal field tensor at the $i$th cell. The cosmic-web classification at the $i$th cell can be compactly expressed as
@@
where $h$ is the Heaviside step function and $\lambda_{\text{th}}$ is a free parameter. In this work, we set $\lambda_{\text{th}}=0.05$ (see Forero–Romero et al. 2009 for the introduction of a threshold in the cosmic-web classification). For recent studies regarding the choice of $\lambda_{\text{th}}$, see Coloma-Nadal et al. (2024) and Olex et al. (2025). $W_{i}$ counts the number of eigenvalues above the threshold; the four resulting values of 0, 1, 2, and 3 correspond to voids, sheets, filaments, and knots, respectively. This cosmic-web definition is problematic if introduced in a forward model that aims to be differentiable with respect to $\boldsymbol{\nu}$. This is because the cosmic-web classification mesh, $\boldsymbol{W,}$ depends on $\boldsymbol{\nu}$ through the potential of $\boldsymbol{\delta}=\mathcal{M}_{\Psi}\circ\mathcal{M}_{\delta}(\boldsymbol{\nu})$, but it does so in a manifestly non-differentiable way due to the presence of the Heaviside step function. As a consequence, the gradients $\partial W_{i}/\partial q_{j}$ are ill-defined, which precludes the use of gradient-based sampling methods. To address this, we relied within the BRIDGE code on a fuzzy cosmic-web classification defined through the sigmoid weights, with
$\boxed{w^{(n)}_{i}=\sigma[s(\lambda^{(n)}_{i}-\lambda_{\text{th}})]}\,$, (13)
@@
where $\sigma(x)=(1+\mathrm{e}^{-x})^{-1}$, and $s$ is the steepness parameter controlling how sharply $w^{(n)}_{i}$ transitions around $\lambda_{\text{th}}$. Because $\sigma$ is strictly increasing and the eigenvalues are sorted, the weights inherit the order $w^{(1)}>w^{(2)}>w^{(3)}$. We converted these cumulative exceedance probabilities into mutually exclusive ‘fuzzy’ memberships for knots (K), filaments (F), sheets (S), and voids (V):
$p^{(\text{K})}_{i}$ $=w^{(3)}_{i}\$, (14)
$p^{(\text{F})}_{i}$ $=w^{(2)}_{i}-w^{(3)}_{i}\$, (15)
$p^{(\text{S})}_{i}$ $=w^{(1)}_{i}-w^{(2)}_{i}\$, (16)
$p^{(\text{V})}_{i}$ $=1-w^{(1)}_{i}\$, (17)
@@
so that they represent the probabilities of having exactly three, two, one, or zero eigenvalues above the threshold, and, by construction, $p^{(\mathrm{V})}_{i}+p^{(\mathrm{S})}_{i}+p^{(\mathrm{F})}_{i}+p^{(\mathrm{K})}_{i}=1$. After introducing this smooth classification scheme, the forward model remained differentiable. Regarding the steepness parameter $s$, a larger value makes the fuzzy classification closer to the original classification, but it also risks causing numerical instabilities in the gradients. We therefore chose the largest $s$ that preserves stable gradient computations. In Fig. 2, we show a comparison of the ‘sharp’ and smooth cosmic-web classification.
@@
We now consider a cosmic-web-dependent bias model where each of the bias parameters from the local-bias model depend on the cosmic-web classification of the voxel. That is, given a bias parameter of $b\in\{\alpha,\rho,\epsilon,\rho^{\prime},\epsilon^{\prime}\}$, it now takes different values in each of the cosmic-web regions: $b^{(\text{K})},b^{(\text{F})},b^{(\text{S})},b^{(\text{V})}$. To ensure differentiability, we used the fuzzy cosmic-web classification, and the bias parameter at voxel $i$ takes the value
$\boxed{b_{i}=b^{(\text{K})}p^{(\text{K})}_{i}+b^{(\text{F})}p^{(\text{F})}_{i}+b^{(\text{S})}p^{(\text{S})}_{i}+b^{(\text{V})}p^{(\text{V})}_{i}}\$. (18)
2.9 Redshift-space-distortion modelling
@@
Redshift-space distortions (RSDs) arise when the observed position of a tracer deviates from its Hubble-flow distance due to peculiar velocities, making them sensitive probes of the growth rate of structure formation (see Kaiser, 1987; Hamilton, 1998). Redshift-space distortions are governed by two distinct types of peculiar motions: coherent large-scale flows, which arise from gravitational infall towards overdense regions, and small-scale random motions within quasi-virialised structures. For a variety of ways of modelling coherent flows, we refer the reader to Kitaura et al. (2012a) and references therein (see also Wang et al., 2012). There are several strategies for incorporating RSDs within a Bayesian framework (we chose the third), which we list below.
@@
    Iterative sampling of the redshift-to-real-space mapping. This approach iteratively reconstructs the real-space distribution from redshift-space observations using Gibbs sampling (Kitaura & Enßlin, 2008; Kitaura et al., 2012c, 2016a). It is inspired by earlier iterative RSD correction methods (see Yahil et al., 1991; Monaco & Efstathiou, 1999).
@@
    Forward modelling of RSDs at the tracer level. Originally introduced by Kitaura et al. (2012b), this method evolves tracers from sampled initial conditions to redshift space, where their final positions are compared with observed galaxy distributions to constrain the primordial density field. Small-scale virial motions can be addressed in two ways: either by collapsing elliptical groups of galaxies – as proposed by Tegmark et al. (2004) – and modelling only coherent flows, as in the original study, or by retaining fingers-of-God features and modelling them with a stochastic velocity dispersion component, as done by Heß et al. (2013).
@@
    Forward modelling of RSDs at the dark-matter-field level. In this work, we followed the approach of mapping the matter field directly to redshift space in a fully differentiable Bayesian inference framework. We set the observer at the center of the cubical volume and computed the RSD effects accordingly along lines of sight. See Bos et al. (2019) for a first implementation of such a method and calculation details.
We note that the modelling of RSDs for biased tracers involving non-linear transformations – such as cosmic voids and Lyman-$\alpha$ forests – requires the inclusion of velocity bias. For the Lyman-$\alpha$ forest, RSD affects the optical depth, $\tau$, but the observable is the transmitted flux $F=\exp(-\tau)$, introducing a non-trivial mapping. Accurate modelling in these cases must account for velocity-bias effects (McDonald et al. (2000); Seljak (2012); Sinigaglia et al. (2022, 2024b)). Similarly, cosmic voids are identified from the galaxy distribution in redshift space, which effectively applies a non-linear and non-local transformation (Chuang et al. (2017)). For the time being, we restrict the implementation tests within the BRIDGE code to unbiased peculiar motions ($b_{v}=1$); i.e. the dark-matter-density contrast in redshift space is given by
$\boxed{\delta(\boldsymbol{s})=\delta\left(\boldsymbol{r}+b_{v}\frac{\boldsymbol{v}\cdot\boldsymbol{r}}{aHr^{2}}\boldsymbol{r}\right)}\,$, (19)
with $\boldsymbol{s}$ and $\boldsymbol{r}$ being the redshift- and real-space coordinates, respectively; $\boldsymbol{v}$ the peculiar velocity; $a$ the scale factor; and $H$ the Hubble constant for a given Hubble distance.
@@ 3 Validation of field-level inference with fuzzy cosmic-web bias
To validate the BRIDGE code, we considered three numerical tests with different resolutions and bias models: TEST1, TEST2, and TEST3 (see Table 1). Results of TEST2 and TEST3 are presented in Appendix C. The numerical tests consist of the following steps.
@@
    Generate a ground-truth number-count catalogue of biased tracers from an initial Gaussian density field with the same forward model employed during inference. All runs are performed to redshift $z=0$ with a particular random seed and specific bias parameters as listed in Tables 2 and 3.
@@
    Perform field-level inference of the posterior describing the white-noise and bias parameters for TEST1 and TEST2, applying the BRIDGE code.
@@
    Perform quality assessment of the reconstructions by evaluating the recovery of both the initial and final conditions using two-point and three-point statistical analyses.
Regarding the hyperparameters of the bias parameter prior of Eq. 4, the power-law ($\alpha$) and negative-binomial ($\beta$) hyperparameters are $\mu_{\alpha}=0.27,\sigma_{\alpha}=0.30,\mu_{\beta}=2.12,and\sigma_{\beta}=0.20$.
Table 1: Setting the different numerical cases. [^2] In all cases, we used a mesh of $128^{3}$ voxels with cubical volumes of distinct side lengths, $L,$ and resolutions, ${\Delta}L=L/N$. We consider the Nyquist frequency $k_{\rm nyq}=\pi/\Delta L$ and the non-isotropic Nyquist frequency $k^{\rm non-iso}_{\rm nyq}=\sqrt{3}\,k_{\rm nyq}$ of a cubical Fourier-space mesh. The output redshift is $z=0$. There are 16 combinations of $\Phi$- and $\delta$-web regions: $j,k\in[1,2,3,4]$.
@@
$\Delta L$ $h^{-1}\,\mathrm{Mpc}$] 10 5 8
$L$ $h^{-1}\,\mathrm{Mpc}$] 1280 640 1024
$k_{\rm nyq}$ $h\,\mathrm{Mpc}^{-1}$] 0.31 0.63 0.39
$k^{\rm non-iso}_{\rm nyq}$ $h\,\mathrm{Mpc}^{-1}$] 0.54 1.10 0.68
non-local bias $\Phi$-webj $\Phi$-webj $\Phi\delta$-webjk
@@
Table 2: Values of bias parameters by cosmic-web region of the reference for the TEST1 and TEST2 cases (8 bias parameters).
@@
Table 3: Values of bias parameters by combined cosmic-web regions for the TEST3 case (64 bias parameters). [^3] We show sub-divisions of $\delta$-web types within each $\Phi$-web classification (see Appendix B and Coloma-Nadal et al. 2024).
@@ 3.1 Motivation of the numerical tests
The numerical tests were chosen to demonstrate a number of scientific relevant cases. We list these below.
@@
  • TEST1 and TEST2: The non-local-bias description used here is based on the $\Phi$-web classification. Balaguera-Antolínez et al. (2018) demonstrated that this model can accurately reproduce key statistical properties of the halo distribution, including one-point PDFs, two-point power spectra at percent-level accuracy, and bispectra with reasonable agreement, generally within 15% (see also Balaguera-Antolínez et al. 2020, Pellejero-Ibañez et al. 2020, and Balaguera-Antolínez et al. 2023, who confirm these results for different halo resolutions). These studies confirm the $\Phi$ web as a physically meaningful and effective framework for modelling non-local bias within a reasonable degree of accuracy through non-parametric approaches. However, for field-level inference applications – such as the one developed here – a parametric bias model is required to enable efficient sampling and marginalisation within a Bayesian framework. The same non-local-bias treatment incorporating an equivalent non-linear-bias prescription to the one considered here (see Eq. 8) was employed by Sinigaglia et al. (2024b, a) to model the Lyman-$\alpha$ forest. Their work demonstrated excellent agreement with high-resolution hydrodynamical simulations, achieving power-spectrum accuracy within 5% up to $k\sim 1$ $h\,\mathrm{Mpc}^{-1}$, along with consistent bispectrum predictions. In both cases, i.e. haloes and Lyman-$\alpha$ forests, the underlying structure-formation model was ALPT. For both the TEST1 and TEST2 cases, we jointly sampled the bias parameters starting with an initial guess randomly sampled from the prior distribution (see Eq. 4).
@@
  • TEST1 focuses on a resolution of 10 $h^{-1}\,\mathrm{Mpc}$, corresponding to a Nyquist frequency of $k\simeq 0.3$ $h\,\mathrm{Mpc}^{-1}$. This resolution surpasses the typical scale used in most cosmological analyses, which are commonly limited to $k<0.2$ $h\,\mathrm{Mpc}^{-1}$(see e.g. Ivanov et al., 2025). Also, the ideal resolution at which BAO reconstruction techniques are applied is slightly lower, i.e.15 $h^{-1}\,\mathrm{Mpc}$ (see e.g. Paillas et al., 2025).
@@
  • TEST2 focuses on a resolution of 5 $h^{-1}\,\mathrm{Mpc}$ with a Nyquist frequency of $k\simeq 0.6$ $h\,\mathrm{Mpc}^{-1}$. This resolution has been reported to be high enough to produce accurate galaxy catalogues combined with a sub-grid model based on ALPT (see Forero Sánchez et al., 2024).
@@
  • TEST3: This case study incorporates the full HICOBIAN non-local bias model, which has been shown to significantly enhance accuracy in both power spectra and bispectra, achieving full compatibility with $N$-body-based catalogues across a wide range of scales. While Coloma-Nadal et al. (2024) demonstrated that this model remains valid at mesh resolutions below 4 $h^{-1}\,\mathrm{Mpc}$, in this study we restricted our analysis to a coarser resolution of 8 $h^{-1}\,\mathrm{Mpc}$. This choice is motivated by the fact that the corresponding Nyquist frequency, $k\simeq 0.4$ $h\,\mathrm{Mpc}^{-1}$, marks the scale beyond which cosmological information is effectively saturated (Hahn & Villaescusa-Navarro, 2021). For computational reasons, we kept the bias parameters fixed in this case, assuming they are known. This assumption is reasonable if the parameters can be extracted from high-fidelity reference catalogues, as was done for the PATCHY BOSS mocks, where the MULTIDARK reference catalogue was designed to reproduce the observed galaxy distribution with halo abundance matching (Rodríguez-Torres et al., 2016). Even more sophisticated reference catalogues are now being developed to reproduce the clustering statistics of the observed DESI catalogues (DESI Collaboration et al., 2025a), from which the parameters of the HICOBIAN model can be accurately extracted (Favole et al., 2025).
We considered $\rho^{\prime}=0$ and $\epsilon^{\prime}=1$ for simplicity, interpreting this by continuity for $\delta>-1$, so the high-density damping factor, $\exp\left[-\left(\frac{1+\delta_{i}}{\rho^{\prime}_{\text{}}}\right)^{-\epsilon^{\prime}_{\text{}}}\right]$, was set to one in all three test cases.